2.- It is known that 20% of the students on campus are smokers. If a total of 10 students are interviewed, to find out their situation regarding tobacco, determine the probability that at least one of the interviewees smokes.

Distribution = ____________________
Answer #1

Solution

Given that ,

p = 0.20

q = 1 - p = 1 - 0.20 = 0.80

n = 10

Using binomial probability formula ,

P(X = x) = (n C x) * p x * (1 - p)n - x

P(X 1  ) = 1 - P( x <1)

= 1 - P(X = 0)

= 1 - (10 C 0) * 0.20 0 * (0.8)10

=1-0.1074

probability=0.8926